OpenStudy (anonymous):

if the function f: x ---> x^2 - x has a domain D= {-1,0,1} find the range .

OpenStudy (anonymous):

Substitute in the values for x to find the range: ie y = (-1)^2 - (-1) y = 1+1 y = 2 so 2 is one of the numbers in the range